OpenStudy (anonymous):

Determine which equation below has a solution of x = 2. A. 5x + 8 – 12x = –6x + 2x – 3 B. 6x + 4 – 12x = –10x + 18x – 3 C. –4x – 6 + x = –2x + x – 10 D. –3x – 8 + x = –4x + x – 10

OpenStudy (camerondoherty):

Well a simpler way than solving each one 1 by one for x you substitue x for 2 and see which one is true

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

u live in michigan ? and the answer is b !

OpenStudy (camerondoherty):

So for A: \[5(2)+ 8 – 12(2) = –6(2)+ 2(2) – 3 \] \[10+8-24=-12+4-3\] \[18+-24=-11\] \[6=-11\] So you do that with all of the equations and see which one works as x=2. I already did A for you, so now you need to do this for B, C, and D

OpenStudy (camerondoherty):

Good Job!